# RAP-SAM [Uncaptioned image]: Towards Real-Time All-Purpose Segment Anything

### v1

@@ 4.1 RAP-SAM Architecture
Table 3: Real-Time All-Purpose Segmentation benchmark. Our proposed RAP-SAM achieves the best accuracy and speed trade-off on three tasks.
@@
Mask2Former [10] R18 35.6 77.5 39.0 30.3 54.7 38.6 89.8G 18.6M 31.2
Mask2Former [10] R50 42.9 79.8 47.6 35.6 58.0 42.1 153G 45.2M 26.6
MaskFormer [11] R18 31.0 76.0 33.2 27.7 55.6 34.1 79.7G 18.5M 38.0
MaskFormer [11] R50 37.4 77.3 41.2 31.8 58.8 40.0 143.0G 45.2M 34.3
MaskFormer [11] TopFormer [92] 31.6 75.9 33.5 28.6 56.1 38.2 60.0G 12.8M 29.9
kMaX-DeepLab [85] R18 27.8 71.0 30.3 24.1 16.9 23.1 87.1G 18.7M 15.0
kMaX-DeepLab [85] R50 36.9 75.7 42.7 28.4 20.1 26.4 280.0G 51.4M 14.9
YOSO [25] R18 31.6 76.8 36.4 24.4 45.7 31.4 57.3G 18.7M 41.0
YOSO [25] R50 37.3 76.9 42.7 29.2 49.2 40.7 119.0G 45.1M 36.3
YOSO [25] TopFormer [92] 31.0 75.9 33.0 26.9 45.1 27.9 36.8G 12.9M 31.1
RAP-SAM R18 39.9 78.6 43.3 34.8 52.7 38.7 60.5G 22.8M 40.3
RAP-SAM R50 46.9 80.8 51.6 39.8 57.9 46.2 123.0G 47.2M 35.1
RAP-SAM TopFormer [92] 34.6 77.0 37.8 29.8 53.3 41.7 40.1G 15.0M 30.7
Light-Weight Decoupled Adapter. After the shared decoder, we also add two lightweight adapters, $A_{obj}$ and $A_{prompt}$ to adapt the shared decoder’s knowledge better. In particular, we use the asymmetric design. $A_{obj}$ uses the same dynamic convolution design to refine the object query further, while $A_{prompt}$ uses pixel-wise cross-attention design. Our key insights are: (1) For panoptic segmentation and video segmentation, the interaction of the scene feature is necessary, but not for interactive segmentation. (2) despite the decoder being shared, the details are still missing for interactive segmentation due to the pooling effect in Equ. 1, where we even find inferior results after adopting a pooling-based adapter. More detailed adapter design can be found in Sec. 5.2.
4.2 Training and Inference
Joint Image and Video Segmentation Co-training. Our goal is to train all segmentation tasks only once jointly. All training targets are one entity label and mask for all three different cases. The entity can be thing, stuff, class-agnostic masks, and their corresponding labels. Note that the instance masks with the same ID $d$ form the tube masks. During training, we apply Hungarian matching between the predicted and ground-truth entity masks to assign object queries to video/image entities and then supervise their predicted masks and classification. The classifier is replaced by CLIP text embedding to avoid cross-dataset taxonomy conflicts.
Following previous works [10, 46], the final loss function is given as $L=\lambda_{cls}L_{m\_cls}+\lambda_{ce}L_{m\_ce}+\lambda_{dice}L_{m\_dice}$. Here, $L_{m\_cls}$ is the Cross-Entropy (CE) loss for mask classification, and $L_{m\_ce}$ and $L_{m\_dice}$ are mask Cross-Entropy (CE) loss and Dice loss [62, 73] for segmentation, respectively. By default, we set $\lambda_{cls}=2,\lambda_{ce}=5,\lambda_{dice}=5$.
Inference. we follow the same inference procedure of Mask2Former [10] for image segmentation. We merge the things and stuff for PS according to the sorted scores. The scores are generated by CLIP text embedding. For video instance segmentation, we use query matching rather than introducing extra tracking components to generate instance ID, following previous work [47, 26, 46]. We adopt a near-online method for inference. For interactive segmentation tasks, we follow the origin SAM [34], by providing box and point prompts and obtaining the binary masks. All the parameters are shared across three different tasks.
Extension and Application. Since our model can perform various segmentation tasks, RAP-SAM can be extended to more datasets and tasks, including semantic segmentation on ADE20k [98] or video panoptic segmentation on VIPSeg [61]. Moreover, our method can support prompt-driven video object segmentation by combining interactive segmentation with a video instance segmentation framework. We put the details of the extension and application for RAP-SAM in the supplementary.
5 Experiments
Datasets. Our benchmark uses COCO [50] and YouTube-VIS 2019 [81] datasets for benchmarking. COCO-SAM is constructed by using ground truth boxes and masks as visual prompt inputs, following the SAM-1B dataset [34]. In addition, to verify the effectiveness and generality of RAP-SAM, we also use other datasets, including ADE-20K [98] and VIP-Seg dataset [61] in Sec. 5.1.
Evaluation Metrics and Devices. For image segmentation, we adopt panoptic quality (PQ), which can be further decomposed into the segmentation quality (SQ) and the recognition quality (RQ). For the VIS task, mAP is adopted. We further report the semantic segmentation results (mIoU) and video panoptic segmentation results (STQ [76] and VPQ [31]). We adopt one A100 GPU for speed testing.
Re-implemented Baselines. Since no previous works explore joint co-training for image, video, and interactive segmentation in one framework. We re-implement several representative baselines using the same codebase, including non-real-time models (Mask2Former [10], kMaX-DeepLab [85]) for reference. In particular, we benchmark over ten different methods, including different backbones and decoders.
Implementation Details. We implement our models and all other baselines in PyTorch [65] with the MMDetection toolbox [5]. We use the distributed training framework with 16 A100 GPUs. Each mini-batch has two images per GPU. In particular, we adopt pseudo video training on COCO by moving image masks with random directions. All the models are trained with 24 epochs. For data augmentation, we adopt large-scale jitter as previous works [10] to build strong baselines. For all models, we adopt training steps and optimizers. Refer to the supplementary for more details.
@@ 5.1 Main Results
Our Benchmark Results. We list our benchmark results in Tab. 3. We benchmark recent state-of-the-art methods using the same training and test settings. From the table, our proposed RAP-SAM achieves the best speed and accuracy trade-off on all three different visual segmentation tasks, under the various backbones. Mask2Former [10] achieves similar or partially stronger results than our method. However, their speed is still limited, which is 7-8 FPS slower than our methods. YOSO [25] runs as fast as our method. However, the performance gap is still significant. Thus, our proposed RAP-SAM is a simple yet effective baseline for real-time and all-purpose settings.
Promptable Results Compared with SAM. In Tab. 4, we verify our trained model as a promptable segmenter, like SAM. We compare our method with SAM-Huge, which is much larger. Despite there being a gap, our method can run fast and with much less GFlops.
Comparison with Specific Design Models on VIP-Seg. We also verify the effectiveness of RAP-SAM on a more challenging video segmentation task, video panoptic segmentation on the VIP-Seg dataset. Compared with recent works [46, 61], RAP-SAM also archives the best speed and accuracy trade-off.
Table 4: Comparison with SAM [34] as a promptable segmenter on various detectors and instance segmenters.
Method Detectors mAP AP50 AP75 Parameters Flops
SAM-Huge Mask-RCNN (R50) 35.6 54.9 38.4 641M 3000G
Ours Mask-RCNN (R50) 26.7 43.0 27.9 47M 188G
SAM-Huge Mask-RCNN (X-101-64x4d) 39.0 60.5 42.1 641M 3000G
Ours Mask-RCNN (X-101-64x4d) 28.5 46.4 29.9 47M 188G
@@
Table 5: Comparison with real-time video segmentation method on VIP-Seg validation set.
Method backbone $\rm VPQ^{1}$ $\rm VPQ^{2}$ $\rm VPQ^{4}$ $\rm VPQ^{6}$ VPQ STQ FPS
Clip-PanoFCN [61] R50 24.3 23.5 22.4 21.6 22.9 31.5 8
Video K-Net [47] R50 29.5 26.5 24.5 23.7 26.1 33.1 10
Tube-Link [46] STDCv1 32.1 31.3 30.1 29.1 30.6 32.0 14
Tube-Link [46] STDCv2 33.2 31.8 30.6 29.6 31.4 32.8 12
Ours R18 34.2 32.1 30.8 30.2 32.5 33.7 30
Table 6: Comparison with real-time panoptic segmentation method on ADE-20k validation set.
Method Backbone PQ $\mathrm{PQ^{th}}$ $\mathrm{PQ^{st}}$ Parameters Flops
PanSegFormer [49] R50 36.4 35.3 38.6 51M 214G
MaskFormer [11] R50 34.7 32.2 39.7 45.2M 181G
Mask2Former [10] R50 39.7 39.0 40.9 45.2M 226G
YOSO [25] R50 38.0 37.3 39.4 45.1M 176G
RAP-SAM R50 38.3 36.6 39.7 47.2M 179G
Comparison with Specific Design Models on ADE-20k. We further transfer RAP-SAM on ADE-20k datasets. Since our method is real-time model, we only list several representative works for reference. Again, compared with recent work [25], our method still achieves stronger results.
5.2 Ablation Study and Visual Analysis
Ablation on Shared Decoder Design. In Tab. 7, we find using simple pooling-based dynamic convolution performs well under real-time settings, where we adopt R18 as the backbone and keep the adapter unchanged. In particular, there is no difference in panoptic segmentation results and a slight gap for interactive segmentation. However, per-pixel cross-attention introduces extra computation costs.
Ablation on Meta-Architecture Design. In Tab. 8, we further explore the meta-architecture design, as shown in Fig. 3. We use R50 as the backbone and keep the adapter unchanged. From the table, we find using a shared decoder architecture (in Tab. 7) achieves the best parameter and performance trade-off. These findings may conflict with recent multi-task works [42] since the decoupled decoders may have more capacity. We argue that since our feature extractors are weaker than those works and the representation power is far from the heavier backbone such as Swin [52] or ViT-large [13], adding more capacity in the decoder may not boost the final performance.
Table 7: Ablation on shared decoder design.
Setting COCO-Panoptic COCO-SAM FPS
Per-Pixel Cross-Attention [10] 45.0 55.3 28.0
Pooling + Dynamic Convolution [68] 43.8 55.7 36.2
Pooling + Dynamic Convolution with Gate [93] 44.6 56.7 34.5
Table 8: Ablation on meta-architecture.
@@ 5.2 Ablation Study and Visual Analysis
Table 9: Ablation on adapter design. CA: cross-attention. DC: dynamic convolution.
@@
Table 10: Ablation on joint co-training. (a), COCO-Panoptic. (b), YT-VIS 2019. (c), COCO-SAM.
Setting COCO-Panoptic YT-VIS 2019 COCO-SAM
a 36.6 - -
b - 21.5 -
a + b 36.2 36.0 50.7
a + b + c 35.7 35.3 50.9
Figure 4: The visualization results of our model on the YouTube-VIS 2019 dataset and the COCO dataset. The first three rows visualize five frames of inputs. The same instances are shown with the same color. The fourth row shows the interactive segmentation results with a single-point prompt (shown in green color). The last row shows the panoptic segmentation results.
Ablation on Adapter Design. Then, we explore the adapter design in Tab. 9. In particular, we use a pre-trained model without adapters for initialization to see the effect of different adapters. In particular, we find using asymmetric adapters works well for balanced results for object queries and prompt queries, since the goals of the two queries are different. The former needs the spatio-temporal and scene-level context, while the latter only cares about the region context with the input position as the guidance.
@@
Effectiveness of Joint Co-training. We further explore the effect of each dataset in Tab. 10. We find joint co-training with image and video data leads to better performance for video instance segmentation. Adding COCO-SAM training leads to a few performance drops. This is because the learning targets are different for object queries and prompt queries.
@@
Visualization. In Fig. 4, we give the three different visualization results on the YouTube-VIS 2019 dataset and the COCO dataset. We generate these visualization results using only a single model thanks to the joint co-training strategy. The first three rows in Fig. 4 show the video instance segmentation. The fourth row shows the interactive segmentation results with a single-point prompt (shown in green color), and the last row shows the panoptic results.

### v2

@@ 3.1 RMP-SAM Architecture
Light-Weight Decoupled Adapter. After the shared decoder, we also add two lightweight adapters, $A_{obj}$ and $A_{prompt}$ to adapt the shared decoder’s knowledge better. In particular, we use the asymmetric design. $A_{obj}$ uses the same dynamic convolution design to refine the object query further, while $A_{prompt}$ uses pixel-wise cross-attention design. Our key insights are: (1) The interaction between the whole scene feature and object queries is necessary for panoptic and video segmentation but not for interactive segmentation. The visual prompts provide strong positional information, and they pay more attention to local details features. (2) Despite the decoder being shared, the details are still missing for interactive segmentation due to the pooling effect in Equ. 1, where we even find inferior results after adopting a pooling-based adapter. More detailed adapter design can be found in Sec. 4.2.
3.2 Training, Inference, and Extension
Joint Image and Video Segmentation Co-training. Our goal is to train all segmentation tasks only once jointly. All training targets are one entity label and mask for all three cases. The entity can be thing, stuff, class-agnostic masks, and corresponding labels. Note that the instance masks with the same ID $d$ form the tube masks. During training, we apply Hungarian matching between the predicted and ground-truth entity masks to assign object queries to video/image entities and then supervise their predicted masks and classification. To unify label taxonomies across datasets, we replace the learnable classifier with CLIP text embeddings following ViLD (Gu et al., 2021).
Following previous works (Cheng et al., 2022; Li et al., 2023e), the final loss function is given as $L=\lambda_{cls}L_{m\_cls}+\lambda_{ce}L_{m\_ce}+\lambda_{dice}L_{m\_dice}$. Here, $L_{m\_cls}$ is the Cross-Entropy (CE) loss for mask classification, and $L_{m\_ce}$ and $L_{m\_dice}$ are mask Cross-Entropy (CE) loss and Dice loss (Milletari et al., 2016; Wang et al., 2020) for segmentation, respectively. By default, we set $\lambda_{cls}=2,\lambda_{ce}=5,\lambda_{dice}=5$.
Inference. We follow the same inference procedure of Mask2Former (Cheng et al., 2022) for image segmentation. In panoptic segmentation, we merge the prediction of things and stuff according to the sorted scores. For video instance segmentation, we use query matching rather than introducing extra tracking components to generate instance ID, following previous work (Li et al., 2022d; Huang et al., 2022; Li et al., 2023e). We adopt a near-online method for inference. We follow the original SAM (Kirillov et al., 2023) for interactive segmentation tasks by providing box and point prompts and obtaining the binary masks. All the parameters are shared across three different tasks.
Extension and Application. Since our model can perform various segmentation tasks, Our method can be extended to more datasets and tasks, including semantic segmentation on ADE20k (Zhou et al., 2017) or video panoptic segmentation on VIPSeg (Miao et al., 2022). Moreover, our method can support prompt-driven video object segmentation by combining interactive segmentation with a video instance segmentation framework. We put the details of the extension and application for RMP-SAM in the appendix (Sec. 7.1).
4 Experiments
Datasets. Our benchmark uses COCO (Lin et al., 2014) and YouTube-VIS 2019 (Yang et al., 2019) datasets for benchmarking. COCO-SAM is constructed using ground truth boxes and points randomly sampled from ground truth masks as visual prompt inputs. In addition, to verify the effectiveness and generality of RMP-SAM, we also use other datasets, including ADE-20K (Zhou et al., 2017) and VIP-Seg dataset (Miao et al., 2022) in Sec. 4.1.
Evaluation Metrics and Devices. For image segmentation, we adopt panoptic quality (PQ), which can be further decomposed into the segmentation quality (SQ) and the recognition quality (RQ). For the VIS task, we use mAP as our primary metric. We further report the semantic segmentation results (mIoU) and video panoptic segmentation results (STQ (Weber et al., 2021) and VPQ (Kim et al., 2020)). Speed testing is conducted fairly on one A100 GPU for all models.
Re-implemented Baselines. Since no previous works explore joint co-training for image, video, and interactive segmentation in one model. We re-implement several representative baselines using the same codebase, including non-real-time models (Mask2Former (Cheng et al., 2022), kMaX-DeepLab (Yu et al., 2022)) for reference. We benchmark multiple different methods, including different backbones and decoders.
Implementation Details. We implement our models and all other baselines in PyTorch (Paszke et al., 2019). We use the distributed training framework with 16 A100 GPUs. Each mini-batch has two images per GPU, and each batch contains one data type. In particular, we adopt pseudo video training on COCO by moving image masks with random directions. All the models are trained with 12 epochs. For data augmentation, we adopt large-scale jitter as previous works (Cheng et al., 2022) to build strong baselines. For all models, we adopt the same training steps and optimizers. Refer to the appendix (Sec. 7.1) for more details.
Table 3: Real-Time Multi-Purpose Segmentation benchmark. Our proposed RMP-SAM achieves the best accuracy and speed trade-off on three tasks.
@@ 4 Experiments
Mask2Former R18 35.6 77.5 39.0 30.3 54.7 38.6 89.8G 18.6M 31.2
MaskFormer R18 31.0 76.0 33.2 27.7 55.6 34.1 79.7G 18.5M 38.0
kMaX-DeepLab R18 27.8 71.0 30.3 24.1 16.9 23.1 87.1G 18.7M 15.0
YOSO R18 31.6 76.8 36.4 24.4 45.7 31.4 57.3G 18.7M 41.0
RMP-SAM(Ours) R18 39.9 78.6 43.3 34.8 52.7 38.7 60.5G 22.8M 40.3
Mask2Former R50 42.9 79.8 47.6 35.6 58.0 42.1 153G 45.2M 26.6
MaskFormer R50 37.4 77.3 41.2 31.8 58.8 40.0 143.0G 45.2M 34.3
kMaX-DeepLab R50 36.9 75.7 42.7 28.4 20.1 26.4 280.0G 51.4M 14.9
YOSO R50 37.3 76.9 42.7 29.2 49.2 40.7 119.0G 45.1M 36.3
RMP-SAM(Ours) R50 46.9 80.8 51.6 39.8 57.9 46.2 123.0G 47.2M 35.1
MaskFormer TopFormer 31.6 75.9 33.5 28.6 56.1 38.2 60.0G 12.8M 29.9
YOSO TopFormer 31.0 75.9 33.0 26.9 45.1 27.9 36.8G 12.9M 31.1
RMP-SAM(Ours) TopFormer 34.6 77.0 37.8 29.8 53.3 41.7 40.1G 15.0M 30.7
@@ 4.1 Main Results
Our Benchmark Results. We list our benchmark results in Tab. 3. We benchmark recent state-of-the-art methods using the same training and test settings. From the table, our proposed RMP-SAM achieves the best speed and accuracy trade-off on all three visual segmentation tasks under the various backbones. Mask2Former (Cheng et al., 2022) achieves similar or partially stronger results than our method. However, their speed is still limited, and they are 7-8 FPS slower than our methods. YOSO (Hu et al., 2023) runs as fast as our method. However, the performance gap is still significant. Thus, our proposed RMP-SAM is a simple yet effective baseline for real-time and Multi-Purpose settings.
Compare with SAM-like Methods. Tab. 4 provides a detailed comparison between our models and previous SAM-like models. We report the results of mask AP of the COCO validation set, the parameter count, and FLOPs of each model. We first generate bounding boxes for each image using Mask R-CNN with ResNet50. Then, we send these bounding boxes to SAM-like models as box prompts. Our method is co-trained using SAM data and demonstrates even better results than most previous efficient SAM-like methods. It has strength in computational efficiency, with an additional function for video instance segmentation and panoptic segmentation.
Comparison with Specific Design Models on VIP-Seg. In Tab. 5(b)(a), we verify the effectiveness of RMP-SAM on a more challenging video segmentation task, video panoptic segmentation (VPS) on the VIP-Seg dataset. Compared with recent works (Li et al., 2023e; Miao et al., 2022), RMP-SAM also archives the best speed and accuracy trade-off, despite RMP-SAM is not specifically designed for VPS.
Table 4: Comparison with SAM-like methods on COCO instance segmentation dataset. We first generate bounding boxes using Faster R-CNN, which are used as visual prompts for segmentation. We report the results of mask AP of the COCO validation set, the parameters, and the FLOPs.
Method mAP #Param. FLOPs(G) PS VIS Interactive
SAM-H (Kirillov et al., 2023) 35.6 641M 3000
TinySAM (Shu et al., 2023) 32.5 10M 42
MobileSAM (Zhang et al., 2023a) 32.1 10M 42
EdgeSAM (Zhou et al., 2023) 31.9 10M 22
EfficientSAM (Xiong et al., 2023) 32.4 10M -
FastSAM (Zhao et al., 2023) 34.3 68M 344
RMP-SAM 33.2 47M 35
@@
Table 5: Compared with specifically designed segmentation models.
(a) Comparison with video segmentation method on VIP-Seg validation set.
Method backbone VPQ STQ FPS
Clip-PanoFCN R50 22.9 31.5 8
Video K-Net R50 26.1 33.1 10
Tube-Link STDCv1 30.6 32.0 14
Tube-Link STDCv2 31.4 32.8 12
Ours R18 32.5 33.7 30
(b) Comparison with panoptic segmentation method on ADE-20k validation set.
Method backbone PQ Parameters FLOPs
PanSegFormer R50 36.4 51M 214G
MaskFormer R50 34.7 45.2M 181G
Mask2Former R50 39.7 45.2M 226G
YOSO R50 38.0 45.1M 176G
RMP-SAM R50 38.3 47.2M 179G
Figure 4: The visualization results of YouTube-VIS 2019 and COCO datasets. The first two rows visualize five frames of inputs. The same instances are in the same color. The third row shows the interactive segmentation results with a single-point prompt (green color). The last row shows the panoptic segmentation results.
Table 6: Ablation study on RMP-SAM’s component design.
(a) Ablation on meta-architecture.
@@
(b) Ablation on joint co-training. a: COCO-Panoptic. b : YT-VIS 2019. c: COCO-SAM.
Setting COCO-Panoptic YT-VIS 2019 COCO-SAM
a 36.6 - -
b - 21.5 -
a + b 36.2 36.0 50.7
a + b + c 35.7 35.3 50.9
(c) Ablation on shared decoder design. DC: dynamic convolution. DCG: dynamic convolution with gating.
Settings COCO-Panoptic COCO-SAM
Per-Pixel Cross-Attention 45.0 55.3
Pooling + DC 43.8 55.7
Pooling + DCG 44.6 56.7
(d) Ablation on adapter design. CA: cross-attention. DC: dynamic convolution.
@@
Comparison with Specific Design Models on ADE-20k. In Tab. 5(b)(b), we further transfer RMP-SAM on ADE-20k datasets. Since our method is a real-time model, we only list several representative works for reference. Our method still achieves stronger results than recent work (Hu et al., 2023).
4.2 Ablation Study and Visual Analysis
Ablation on Meta-Architecture Design. In Tab. 6(d)(a), we further explore the meta-architecture design, as shown in Fig. 3. We use R50 as the backbone and keep the adapter unchanged. The table shows that a shared decoder architecture achieves the best parameter and performance trade-off. These findings may conflict with recent multi-task work (Li et al., 2022c) since the decoupled decoders may have more capacity. We argue that since our feature extractors are weaker than those works and the representation power is far from the heavier backbone such as Swin (Liu et al., 2021) or ViT-large (Dosovitskiy et al., 2021), adding more capacity in the decoder may not boost the final performance.
@@ 4.2 Ablation Study and Visual Analysis
Effectiveness of Joint Co-training. We further explore the impact of each dataset in Tab. 6(d)(b). We find joint co-training with image and video data leads to better performance for video instance segmentation but reduces the performance of panoptic segmentation. Adding COCO-SAM training leads to a few performance drops. This is because the learning targets are different for object queries and prompt queries.
Ablation on Shared Decoder Design. Tab. 6(d)(c) shows using simple pooling-based dynamic convolution performs well under real-time settings, where we adopt ResNet18 as the backbone and keep the adapter unchanged. In particular, there is no difference in panoptic segmentation results and a slight gap for interactive segmentation. However, per-pixel cross-attention introduces extra computation costs.
Ablation on Adapter Design. Then, we explore the adapter design in Tab. 6(d)(d). In particular, we use a pre-trained model without adapters for initialization to see the effect of different adapters. In particular, we find that using asymmetric adapters works well for balanced results for object queries and prompt queries since the goals of the two queries are different. The former needs the spatio-temporal and scene-level context, while the latter only cares about the region context with the input position as the guidance. Our adapter can also work with other architecture, as shown in the appendix (Sec. 7.2).
@@
Result Visualization. In Fig. 4, we give the three different visualization results on the YouTube-VIS 2019 and COCO datasets. We generate these visualization results using only a single model thanks to the joint co-training strategy. The first three rows in Fig. 4 show the video instance segmentation. The fourth row shows the interactive segmentation results with a single-point prompt (shown in green color), and the last row shows the results of panoptic segmentation.
5 Related Work
Universal Segmentation. Prior research has focused on designing segmentation models tailored for specific tasks. Recent developments in this field (Dosovitskiy et al., 2021; Liu et al., 2021; Carion et al., 2020; Zhang et al., 2021; Cheng et al., 2022; Yu et al., 2022; Li et al., 2021b; 2023c; Gu et al., 2023; Yuan et al., 2022; Li et al., 2022c; 2023d; 2024b; Yuan et al., 2024; Yang et al., 2021b; 2024) have shifted towards employing mask classification architectures, coupled with an end-to-end set prediction objective, to facilitate universal segmentation. This approach has outperformed specialized models (Chen et al., 2019; Kirillov et al., 2019; He et al., 2017; Li et al., 2021a; Huang et al., 2019; Yuan et al., 2020; Zhao et al., 2017; Chen et al., 2018; Li et al., 2022b; 2020a; Chen et al., 2017) across various segmentation tasks, including those involving images, videos, and point clouds (Kim et al., 2022; Li et al., 2022d; 2023e; Cheng et al., 2021a; Xu et al., 2022). Specifically, Mask2Former (Cheng et al., 2022) uses a masked-attention mechanism and surpasses the performance of prior specific image segmentation models. Similarly, Tube-Link (Li et al., 2023e) outperforms earlier models (Li et al., 2022d; b; Liu et al., 2020; Shin et al., 2023; Goel et al., 2021) tailored for specific video segmentation tasks. Semantic-SAM (Li et al., 2023a), built on Mask-DINO (Li et al., 2023b), employs a hybrid query design to facilitate panoptic and interactive segmentation within a single framework. However, despite the strong performance of these works, their complexity, stemming from the heavier encoder or the cascaded decoder designs, poses significant challenges for integration into actual products. Slow processing speeds and substantial parameter overhead primarily hinder these.
Efficient Model Design. This research direction primarily concentrates on the development of efficient CNNs (Howard et al., 2017; Sandler et al., 2018; Howard et al., 2019; Iandola et al., 2016; Zhang et al., 2018; Ma et al., 2018; Han et al., 2020), transformers (Maaz et al., 2022; Pan et al., 2022; Mehta & Rastegari, 2022), and hybrid architectures (Zhang et al., 2023b; Ma et al., 2022; Chen et al., 2022b; Li et al., 2022a; Zhou et al., 2022), aimed at advancing visual representation learning. Most research investigates backbone design within the constraints of real-time processing and parameter efficiency. These studies are orthogonal to our work, which primarily explores the impact of data, task, and decoder design. Additionally, our work employs efficient models as encoders and presents detailed benchmarks and analyses of the encoder’s effects.
Efficient Segmentation. Previous studies (Zhao et al., 2018; Li et al., 2020b; Yu et al., 2018; Hu et al., 2023; Hong et al., 2021; Wan et al., 2023; Yu et al., 2021; Mehta et al., 2018) on efficient segmentation have predominantly concentrated on closed-set and specific domains. In particular, a significant portion of this research (Li et al., 2020b; Hu et al., 2023; Mehta et al., 2019) is dedicated to driving scenarios. Recently, various studies (Zhang et al., 2022; Wan et al., 2023; Zhou et al., 2023) have developed efficient segmentation techniques that facilitate model execution on mobile devices. Mobile SAM (Zhang et al., 2023a) introduces a streamlined encoder distillation method. Fast SAM (Zhao et al., 2023) employs a single-stage instance segmentation framework that directly decodes class-agnostic masks. Recently, multiple studies have been conducted on efficient panoptic segmentation (Sun et al., 2023; Hu et al., 2023; Mohan & Valada, 2021) and rapid video instance segmentation (Yang et al., 2021a; Zhang et al., 2024). However, these real-time segmentation methods are limited to specific tasks. We contend that a multi-purpose model capable of real-time performance would have wide-ranging applications in editing, tracking, and segmentation functionalities within various products.
Interactive Segmentation. Interactive segmentation (Kirillov et al., 2023; Chen et al., 2022a; Sofiiuk et al., 2022; Lin et al., 2020; 2022) distinguishes objects by actively incorporating user inputs. Recently, SAM (Kirillov et al., 2023) has shown remarkable segmentation outcomes by leveraging multiple user interaction inputs and training on SA-1B data. Inspired by SAM, several studies (Zhang et al., 2023a; Zhou et al., 2023) have concentrated on employing a lighter image encoder to adapt SAM for real-time scenarios.